### v3

@@ I Introduction
While the Sibson and Arimoto MI have closed-form expressions, the Augustin–Csiszár MI, Lapidoth–Pfister MI, and $\alpha$-capacity generally do not; therefore, these types of MI and $\alpha$-capacity must be calculated by solving optimization problems. Algorithms for computing the $\alpha$-MI and $\alpha$-capacity have been explored [2, 16, 17, 3, 18, 19, 20]. For the Augustin–Csiszár MI for $\alpha\in(0,1)$, Augustin [3] proposed an iterative algorithm, which Karakos et al. [18] rediscovered as an alternating optimization (AO) algorithm. Recently, Wang et al. [19] proposed an algorithm based on geodesically convex optimization, which is effective for $\alpha\in(1,\infty)$. For the Sibson and Arimoto capacities, Arimoto [2, 16] proposed AO algorithms similar to the Arimoto–Blahut algorithm [21, 22] for calculating the channel capacity $C$. Recently, Kamatsuka et al. [17] proposed additional AO algorithms for computing these capacities and demonstrated that they are equivalent to those suggested by Arimoto. For Augustin–Csiszár capacity, Kamatsuka et al. [20] proposed an AO algorithm. Table I summarizes the $\alpha$-MI, $\alpha$-capacity, and their calculation methods.
@@ II Preliminaries
Karakos et al. [18], Shayevitz and Ofer [23], and Lapidoth and Pfister [5] provided the following variational characterizations of $\alpha$-MI.
Proposition 2 ([18], [23, Thm 1], [5, Lemma 7]).
@@
$=\min_{\tilde{q}_{X,Y}}\{\alpha D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X})$
$\qquad\qquad\qquad+(1-\alpha)I(p_{X},\tilde{q}_{Y\mid X})+D(\tilde{q}_{X}||p_{X})\},$ (12)
@@ IV-C AO Algorithms for Computing $\alpha$-Capacity
Arimoto [2] proposed an AO algorithm for computing the Sibson capacity based on the characterization in (14). Subsequently, Arimoto [16] and Kamatsuka et al. [17] proposed similar AO algorithms, building on the characterizations of Arimoto and Sibson MI in Remark 2 and (15) (see [17, Thm 1 and Table I]). Kamatsuka et al. [20] derived an AO algorithm for computing the Augustin–Csiszár capacity for $\alpha\in(1,\infty)$ based on the characterization in (20). In [20], they also proposed another AO algorithm for the Sibson capacity by modifying the AO algorithm for computing the correct decoding exponent proposed by Jitsumatsu and Oohama [26, Algorithm 1].
@@
The AO algorithm with the above updating formulae is equivalent to the AO algorithm for computing the correct decoding exponent without a cost constraint in [26, Algorithm 2].
@@ IV-D1 AO algorithms for computing $\alpha$-capacity
Arimoto [27, Thm 3] and Kamatsuka et al. [17, Cor 2] demonstrated the global convergence properties of the AO algorithms based on the characterization of the Sibson MI in (14) and (15) as well as the corresponding AO algorithms derived from the characterization of the Arimoto MI. The modified Jitsuamtsu–Oohama algorithm proposed in [20] also has this property (see [26, Thm 2]), while the property of the AO algorithm based on the characterization in (20) for computing the Augustin–Csiszár capacity for $\alpha\in(1,\infty)$ remains undemonstrated.
@@
where $\tilde{q}_{X,Y}^{\text{opt}}$ is the optimal joint distribution that achieves the maximum in (24).
@@
$\leq\frac{\alpha}{\alpha-1}\sum_{x,y}\tilde{q}_{X,Y}^{*}(x,y)\log\frac{\tilde{q}_{X,Y}^{(N+1)}(x,y)}{\tilde{q}_{X,Y}^{(0)}(x,y)}$
$\qquad\qquad\qquad+\sum_{x}\tilde{q}_{X}^{*}(x)\log\frac{\tilde{q}_{X}^{(N+1)}(x)}{\tilde{q}_{X}^{(0)}(x)}$ (49)
$=\frac{\alpha}{\alpha-1}\left(D(\tilde{q}_{X,Y}^{*}||\tilde{q}_{X,Y}^{(0)})-D(\tilde{q}_{X,Y}^{*}||\tilde{q}_{X,Y}^{(N+1)})\right)$
$\qquad\qquad\qquad+D(\tilde{q}_{X}^{*}||\tilde{q}_{X}^{(0)})-D(\tilde{q}_{X}^{*}||\tilde{q}_{X}^{(N+1)})$ (50)
$\leq\frac{\alpha}{\alpha-1}D(\tilde{q}_{X,Y}^{*}||\tilde{q}_{X,Y}^{(0)})+D(\tilde{q}_{X}^{*}||\tilde{q}_{X}^{(0)}).$ (51)
@@ IV-D2 AO algorithm for computing Augustin–Csiszár MI
Karakos et al. [18, Thm 1] proved that the AO algorithm based on (20) and (22) converges to the Augustin–Csiszár MI for $\alpha\in(0,1)$. Similar to the proof of convergence of the Lapidoth–Pfister capacity, we show that the AO algorithm based on (20) and (23) globally converges when $\alpha\in(1,\infty)$. Hence, it suffices to prove the following lemmas:
@@ IV-D3 AO algorithms for computing Lapidoth–Pfister MI
The global convergence property of the AO algorithm based on Corollary 1 follows immediately from the sufficient condition for the global convergence of general AO algorithms [28, Thm 10.5].
@@
We can prove (56) by combining the sufficient condition from [28, Thm 10.5] with the fact that $(q_{X},q_{Y})\mapsto D_{\alpha}(p_{X}p_{Y\mid X}||q_{X}q_{Y})$ is jointly convex for $[1/2,1)\cup(1,\infty)$ [5, Lemma 15]. ∎
@@
where $\tilde{q}_{X}^{\text{opt}}(x):=\sum_{y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)$ and $\tilde{q}_{X,Y}^{\text{opt}}$ is the optimal joint distributions that achieve the minimum in (24).
@@ V Comparison of AO algorithms for $\alpha$-Capacity
It is known that the Sibson, Arimoto, and Augustin–Csiszár capacity are equivalent: $C_{\alpha}^{\text{S}}=C_{\alpha}^{\text{A}}=C_{\alpha}^{\text{C}}$ for $\alpha\in(0,1)\cup(1,\infty)$ [2],[4],[30]. Moreover, the Lapidoth–Pfister capacity $C_{\alpha}^{\text{LP}}$ is equivalent to these capacities for $\alpha\in(1,\infty)$ [30, Thm 4], [31, p.4]. Thus, all AO algorithms obtained so far compute equivalent quantities. This section compares their performance through numerical examples. From this point onward, we refer to an AO algorithm based on the variational characterization using a functional $F_{\alpha}^{\text{X}}(\cdot)$ or $\tilde{F}_{\alpha}^{\text{X}}(\cdot)$ as Algorithm X. Moreover, we refer to the modified AO algorithm of [26, Algorithm 1] as Algorithm JO.
Kamatsuka et al. [17, Cor 1] proved that the AO algorithms S1, S2, and the algorithms based on the variational characterizations of the Arimoto MI are all equivalent under certain conditions on initial distributions. Similarly to [26, Lemma 7 and 8], it can be shown that the Algorithm S1 achieves a greater increase in the objective function per update than Algorithm S3. Thus, we numerically compared the AO algorithms S1, JO, C, and LP for $\alpha\in(1,\infty)$.
@@ Appendix A Proof of Theorem 1
where $(a)$ follows from (12), $(b)$ follows from (3), $(c)$ follows from the minimax theorem (see, e.g., [32, Thm 4.2]) [^3] Notably, $\tilde{F}_{\alpha}^{\text{S3}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y}):=\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})+\mathbb{E}_{X,Y}^{\tilde{q}_{X,Y}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{\tilde{q}_{X}(X)}\right]+\frac{1}{1-\alpha}D(\tilde{q}_{X}||p_{X}),\alpha\in(0,1)$ is concave for $r_{X\mid Y}$ (see, e.g. [28, Sec. 10.3.2]) and convex for $\tilde{q}_{X,Y}$, which can be easily verified., and $(d)$ follows from (62) and the properties of the Kullback–Leibler divergence (see, e.g., [33, Thm 2.6.3]). For $\alpha\in(1,\infty)$, Theorem 1 can be shown by replacing $\min_{\tilde{q}_{X,Y}}$ with $\max_{\tilde{q}_{X,Y}}$ in (63)–(67).
@@ Appendix B Proof of Theorem 2
where $\hat{q}_{Y\mid X}(\cdot|x)$ is defined in (68) and $(a)$ follows from [23, Thm 1], $(b)$ follows from (3), $(c)$ follows from the minimax theorem [^4] Notably, $\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},r_{X\mid Y}):=\mathbb{E}_{X,Y}^{p_{X}\tilde{q}_{Y\mid X}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{p_{X}(X)}\right]+\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X}),\alpha\in(0,1)$ is concave for $r_{X\mid Y}$ and linear (hence, convex) for $\tilde{q}_{Y\mid X}$. , and $(d)$ follows from the properties of the Kullback–Leibler divergence (see, e.g., [33, Thm 2.6.3]). Eq. (20) for $\alpha\in(0,1)$ can be obtained from (1) and (70) by replacing $I(p_{X},\tilde{q}_{Y\mid X})$ with $\min_{q_{Y}}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}q_{Y})$, which was originally proved by Karakos et al. [18]. For $\alpha\in(1,\infty)$, Theorem 2 can be shown by replacing $\min_{\tilde{q}_{Y\mid X}}$ with $\max_{\tilde{q}_{Y\mid X}}$ in Eqs. (70)–(73).
@@ Appendix E Proof of Lemma 2
×$\times\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)\log\frac{p_{X}^{(k+1)}(x)^{\frac{\alpha}{2\alpha-1}}p_{Y\mid X}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{X,Y}^{(k)}(x,y)^{\frac{\alpha-1}{2\alpha-1}}}{\tilde{q}_{X}^{(k+1)}(x)\tilde{q}_{Y\mid X}^{(k+1)}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{Y}^{(k)}(y)^{\frac{\alpha-1}{2\alpha-1}}}$ (96)
@@
$=\frac{1}{1-\alpha}\mathbb{E}_{X}^{\tilde{q}_{X}^{\text{opt}}}\left[D(\tilde{q}_{Y\mid X}^{\text{opt}}(\cdot\mid X)||\tilde{q}_{Y\mid X}(\cdot\mid X))\right]$
$+\frac{\alpha}{\alpha-1}\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}\log\frac{\tilde{q}_{X,Y}^{k+1}(x,y)}{\tilde{q}_{X,Y}^{(k)}(x,y)}$
@@ Appendix G Proof of Lemma 6
where $(f)$ follows from the nonnegativity of the Kullback–Leibler divergence and $\frac{1-2\alpha}{1-\alpha}\leq 0$ for $\alpha\in[1/2,1]$ and $(g)$ follows from the data-processing inequality (also known as monotonicity) of the Kullback–Leibler divergence (see, e.g., [24, Thm 2.17]). ∎

### v5

@@ I Introduction
While the Sibson and Arimoto MI have closed-form expressions, the Augustin–Csiszár MI, Lapidoth–Pfister MI, and $\alpha$-capacity generally do not; therefore, these types of MI and $\alpha$-capacity must be calculated by solving optimization problems. Algorithms for computing the $\alpha$-MI and $\alpha$-capacity have been explored in [2, 16, 17, 3, 18, 19, 20, 21, 22, 23]. For the Augustin–Csiszár MI for $\alpha\in(0,1)$, Augustin [3] proposed an iterative algorithm, which Karakos et al. [18] rediscovered as an alternating optimization (AO) algorithm. Recently, other iterative algorithms have been proposed in [19, 20, 21, 22], which are effective for $\alpha\in(1,\infty)$. For the Sibson and Arimoto capacities, Arimoto [2, 16] proposed AO algorithms similar to the Arimoto–Blahut algorithm [24, 25] for calculating the channel capacity $C$. Recently, Kamatsuka et al. [17] proposed additional AO algorithms for computing these capacities and demonstrated that they are equivalent to those suggested by Arimoto. For Augustin–Csiszár capacity, Kamatsuka et al. [23] proposed an AO algorithm. Table I summarizes the $\alpha$-MI, $\alpha$-capacity, and their calculation methods.
@@ II Preliminaries
Csiszár [4], Karakos et al. [18], Shayevitz [26], and Lapidoth and Pfister [5] provided the following variational characterizations of $\alpha$-MI.
Proposition 2 ([4, p. 34], [18], [26, Thm 1], [5, Lemma 8]).
@@
$=\min_{\tilde{q}_{X,Y}}\{\alpha D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})$
$\qquad\qquad\qquad+(1-\alpha)I(\tilde{q}_{X},\tilde{q}_{Y\mid X})+D(\tilde{q}_{X}||p_{X})\},$ (12)
@@ III Variational Characterizations of $\alpha$-MI
Remark 3.
It is worth mentioning that Poltyrev [29, Thm 3.4] reported the equivalent characterization in (21) (see [13, Lemma 18]).
@@ IV-C AO Algorithms for Computing $\alpha$-Capacity
Arimoto [2] proposed an AO algorithm for computing the Sibson capacity based on the characterization in (14). Subsequently, Arimoto [16] and Kamatsuka et al. [17] proposed similar AO algorithms, building on the characterizations of Arimoto and Sibson MI in Remark 2 and (15) (see [17, Thm 1 and Table I]). Kamatsuka et al. [23] derived an AO algorithm for computing the Augustin–Csiszár capacity for $\alpha\in(1,\infty)$ based on the characterization in (20). In [23], they also proposed another AO algorithm for the Sibson capacity by modifying the AO algorithm for computing the correct decoding exponent proposed by Jitsumatsu and Oohama [30, Algorithm 1].
@@
The AO algorithm with the above updating formulae is equivalent to the AO algorithm for computing the correct decoding exponent without a cost constraint in [30, Algorithm 2].
@@ IV-D1 AO algorithms for computing $\alpha$-capacity
Arimoto [31, Thm 3] and Kamatsuka et al. [17, Cor 2] demonstrated the global convergence properties of the AO algorithms based on the characterization of the Sibson MI in (14) and (15) as well as the corresponding AO algorithms derived from the characterization of the Arimoto MI. The modified Jitsuamtsu–Oohama algorithm proposed in [23] also has this property (see [30, Thm 2]), while the property of the AO algorithm based on the characterization in (20) for computing the Augustin–Csiszár capacity for $\alpha\in(1,\infty)$ remains undemonstrated.
@@
where $\tilde{q}_{X,Y}^{\text{opt}}$ be the optimal distributions that attain the Lapidoth–Pfister capacity $C_{\alpha}^{\text{LP}}=\max_{p_{X}}\max_{\tilde{q}_{X,Y}}\max_{r_{X\mid Y}}\tilde{F}_{\alpha}^{\text{LP}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$.
@@
$\leq\frac{\alpha}{\alpha-1}\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)\log\frac{\tilde{q}_{X,Y}^{(N+1)}(x,y)}{\tilde{q}_{X,Y}^{(0)}(x,y)}$
$\qquad\qquad\qquad+\sum_{x}\tilde{q}_{X}^{\text{opt}}(x)\log\frac{\tilde{q}_{X}^{(N+1)}(x)}{\tilde{q}_{X}^{(0)}(x)}$ (49)
$=\frac{\alpha}{\alpha-1}\left(D(\tilde{q}_{X,Y}^{\text{opt}}||\tilde{q}_{X,Y}^{(0)})-D(\tilde{q}_{X,Y}^{\text{opt}}||\tilde{q}_{X,Y}^{(N+1)})\right)$
$\qquad\qquad\qquad+D(\tilde{q}_{X}^{\text{opt}}||\tilde{q}_{X}^{(0)})-D(\tilde{q}_{X}^{\text{opt}}||\tilde{q}_{X}^{(N+1)})$ (50)
$\leq\frac{\alpha}{\alpha-1}D(\tilde{q}_{X,Y}^{\text{opt}}||\tilde{q}_{X,Y}^{(0)})+D(\tilde{q}_{X}^{\text{opt}}||\tilde{q}_{X}^{(0)}).$ (51)
@@ IV-D2 AO algorithm for computing Augustin–Csiszár MI
Augustin [3, Lemma 34.2], Karakos et al. [18, Thm 1], and Nakiboǧulu [13, Lemma 13 (c)] proved that the AO algorithm based on (20) and (22) converges to the Augustin–Csiszár MI for $\alpha\in(0,1)$. Similar to the proof of convergence of the Lapidoth–Pfister capacity, we show that the AO algorithm based on (20) and (23) globally converges when $\alpha\in(1,\infty)$. Hence, it suffices to prove the following lemmas:
@@ IV-D3 AO algorithms for computing Lapidoth–Pfister MI
The global convergence property of the AO algorithm based on Corollary 1 follows immediately from the sufficient condition for the global convergence of general AO algorithms [32, Thm 10.5].
@@
We can prove (56) by combining the sufficient condition from [32, Thm 10.5] with the fact that $(q_{X},q_{Y})\mapsto D_{\alpha}(p_{X}p_{Y\mid X}||q_{X}q_{Y})$ is jointly convex for $[1/2,1)\cup(1,\infty)$ [5, Lemma 15]. ∎
@@
where $\tilde{q}_{X}^{\text{opt}}(x):=\sum_{y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)$ and $\tilde{q}_{X,Y}^{\text{opt}}$ are the optimal joint distributions that achieve the minimum in (24).
@@ V Comparison of AO algorithms for $\alpha$-Capacity
It is known that the Sibson, Arimoto, and Augustin–Csiszár capacity are equivalent: $C_{\alpha}^{\text{S}}=C_{\alpha}^{\text{A}}=C_{\alpha}^{\text{C}}$ for $\alpha\in(0,1)\cup(1,\infty)$ [2],[4],[34]. Moreover, the Lapidoth–Pfister capacity $C_{\alpha}^{\text{LP}}$ is equivalent to these capacities for $\alpha\in(1,\infty)$ [34, Thm 4], [35, p.4]. Thus, all AO algorithms obtained so far compute equivalent quantities. This section compares their performance through numerical examples. From this point onward, we refer to an AO algorithm based on the variational characterization using a functional $F_{\alpha}^{\text{X}}(\cdot)$ or $\tilde{F}_{\alpha}^{\text{X}}(\cdot)$ as Algorithm X. Moreover, we refer to the modified AO algorithm of [30, Algorithm 1] as Algorithm JO.
Kamatsuka et al. [17, Cor 1] proved that the AO algorithms S1, S2, and the algorithms based on the variational characterizations of the Arimoto MI are all equivalent under certain conditions on initial distributions. Similarly to [30, Lemma 7 and 8], it can be shown that the Algorithm S1 achieves a greater increase in the objective function per update than Algorithm S3. Thus, we numerically compared the AO algorithms S1, JO, C, and LP for $\alpha\in(1,\infty)$.
@@ VI Conclusion
Acknowledgment
The authors would like to thank Barış Nakiboǧlu and anonymous reviewers for their helpful comments.
@@ Appendix A Proof of Theorem 1
where $(a)$ follows from (12), $(b)$ follows from (3), $(c)$ follows from the minimax theorem (see, e.g., [36, Thm 4.2]) [^3] Notably, $\tilde{F}_{\alpha}^{\text{S3}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y}):=\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})+\mathbb{E}_{X,Y}^{\tilde{q}_{X,Y}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{\tilde{q}_{X}(X)}\right]+\frac{1}{1-\alpha}D(\tilde{q}_{X}||p_{X}),\alpha\in(0,1)$ is concave for $r_{X\mid Y}$ (see, e.g. [32, Sec. 10.3.2]) and convex for $\tilde{q}_{X,Y}$, which can be easily verified., and $(d)$ follows from (62) and the properties of the Kullback–Leibler divergence (see, e.g., [37, Thm 2.6.3]). For $\alpha\in(1,\infty)$, Theorem 1 can be shown by replacing $\min_{\tilde{q}_{X,Y}}$ with $\max_{\tilde{q}_{X,Y}}$ in (63)–(67).
@@ Appendix B Proof of Theorem 2
where $\hat{q}_{Y\mid X}(\cdot|x)$ is defined in (68) and $(a)$ follows from [26, Thm 1], $(b)$ follows from (3), $(c)$ follows from the minimax theorem [^4] Notably, $\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},r_{X\mid Y}):=\mathbb{E}_{X,Y}^{p_{X}\tilde{q}_{Y\mid X}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{p_{X}(X)}\right]+\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X}),\alpha\in(0,1)$ is concave for $r_{X\mid Y}$ and linear (hence, convex) for $\tilde{q}_{Y\mid X}$. , and $(d)$ follows from the properties of the Kullback–Leibler divergence (see, e.g., [37, Thm 2.6.3]). Eq. (20) for $\alpha\in(0,1)$ can be obtained from (1) and (70) by replacing $I(p_{X},\tilde{q}_{Y\mid X})$ with $\min_{q_{Y}}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}q_{Y})$, which was originally proved by Karakos et al. [18]. For $\alpha\in(1,\infty)$, Theorem 2 can be shown by replacing $\min_{\tilde{q}_{Y\mid X}}$ with $\max_{\tilde{q}_{Y\mid X}}$ in Eqs. (70)–(73).
@@ Appendix E Proof of Lemma 2
×$\times\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)$
×$\times\log\frac{p_{X}^{(k+1)}(x)^{\frac{\alpha}{2\alpha-1}}p_{Y\mid X}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{X,Y}^{(k)}(x,y)^{\frac{\alpha-1}{2\alpha-1}}}{\tilde{q}_{X}^{(k+1)}(x)\tilde{q}_{Y\mid X}^{(k+1)}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{Y}^{(k)}(y)^{\frac{\alpha-1}{2\alpha-1}}}$ (96)
@@
$=\frac{1}{1-\alpha}\mathbb{E}_{X}^{\tilde{q}_{X}^{\text{opt}}}\left[D(\tilde{q}_{Y\mid X}^{\text{opt}}(\cdot\mid X)||\tilde{q}_{Y\mid X}^{(k)}(\cdot\mid X))\right]$
$+\frac{\alpha}{\alpha-1}\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)\log\frac{\tilde{q}_{X,Y}^{k+1}(x,y)}{\tilde{q}_{X,Y}^{(k)}(x,y)}$
@@ Appendix G Proof of Lemma 6
where $(f)$ follows from the nonnegativity of the Kullback–Leibler divergence and $\frac{1-2\alpha}{1-\alpha}\leq 0$ for $\alpha\in[1/2,1]$ and $(g)$ follows from the data-processing inequality (also known as monotonicity) of the Kullback–Leibler divergence (see, e.g., [27, Thm 2.17]). ∎